# Diffusion Models, Image Super-Resolution And Everything: A Survey

### v2

@@ VII-G Cascaded Image Generation
Saharia et al. [15] presented the concept of cascaded image generation, in which multiple DDPMs are chained together across different scales. This strategy was applied to unconditional and class-conditional generation, cascading a model synthesizing $64\times 64$ images with SR3 models generating $1024\times 1024$ unconditional faces and $256\times 256$ class-conditional natural images. The cascading approach allows several simpler models to be trained simultaneously, improving computational efficiency due to faster training times and reduced parameter counts. Furthermore, they implemented cascading for inference, using more refinement steps at lower and fewer steps at higher resolutions. They found this more efficient than generating SR images directly while achieving satisfactory FID scores. Even though their model’s underperformance compared to BigGAN [110] concerning cascaded generation, the approach still represents an exciting research opportunity.
VIII Conclusion
@@
In image Super-Resolution (SR), integrating Diffusion Models (DMs) has ushered in a new era of innovation. DMs, as exemplified by SR3, have the potential to bridge the gap between technical image quality and human perceptual preferences. Unlike traditional SR methods, which often focus solely on pixel-level accuracy, DMs have the unique ability to generate images that are high in resolution, aesthetically pleasing, and realistic to the human eye. In contrast to previous generative models, they do not suffer typical training convergence issues like mode collapse. Throughout this survey, we have explored the remarkable progress and diverse methodologies that have propelled SR using DMs to the forefront of the field. We began by establishing the foundational principles of DMs, highlighting their relationship to other generative approaches, elucidating their core concepts and their application to image SR. The survey unveiled the myriad of conditioning strategies, from LR image guidance to text embeddings and pre-trained networks, showcasing their crucial role in enhancing image quality. Zero-shot SR, a particularly intriguing paradigm, was also a subject, as well as corruption spaces and image SR-specific topics like color shifting and architectural designs.
In conclusion, this survey unveils the profound impact of DMs on image SR. It provides a comprehensive guide to the current landscape and valuable insights into trends, challenges, and future directions. DMs do not just close a significant portion of the gap in SR technology; they set a new benchmark, pushing the boundaries of image quality and realism. As we continue to explore and refine these models, the future of image SR looks more promising than ever, with potential applications, as discussed in our applications section, that extend far beyond what was previously imagined.
For future research, this survey highlights important issues, such as color shifting, computational costs, efficient sampling, and comparability, while also discussing interesting avenues like image manipulation and cascaded image generation.

### v3

@@ VIII-G Cascaded Image Generation
Saharia et al. [15] presented cascaded image SR, in which multiple DDPMs are chained across different scales. This strategy was applied to unconditional and class-conditional generation, cascading a model synthesizing $64\times 64$ images with SR3 models generating $1024\times 1024$ unconditional faces and $256\times 256$ class-conditional natural images. The cascading approach allows several simpler models to be trained simultaneously, improving computational efficiency due to faster training times and reduced parameter counts. Furthermore, they implemented cascading for inference, using more refinement steps at lower and fewer steps at higher resolutions. They found this more efficient than generating SR images directly. Even though their approach underperforms compared to BigGAN [102] concerning cascaded generation, it still represents an exciting research opportunity.
IX Conclusion
@@
Diffusion Models (DMs) revolutionized image Super-Resolution (SR) by enhancing both technical image quality and human perceptual preferences. While traditional SR often focuses solely on pixel-level accuracy, DMs can generate HR images that are aesthetically pleasing and realistic. Unlike previous generative models, they do not suffer typical convergence issues. This survey explored the progress and diverse methods that have propelled DMs to the forefront of SR. Potential use cases, as discussed in our applications section, extend far beyond what was previously imagined. We introduced their foundational principles and compared them to other generative models. We explored conditioning strategies, from LR image guidance to text embeddings. Zero-shot SR, a particularly intriguing paradigm, was also a subject, as well as corruption spaces and image SR-specific topics like color shifting and architectural designs. In conclusion, the survey provides a comprehensive guide to the current landscape and valuable insights into trends, challenges, and future directions. As we continue to explore and refine these models, the future of image SR looks more promising than ever.